Helsinki University of Technology Publications in Telecommunications Software and Multimedia Teknillisen korkeakoulun tietoliikenneohjelmistojen ja multimedian julkaisuja Espoo 1999 TML-XX

IP network management

Version August 30, 1999

Liisa Uosukainen Tuomas Lilja Lasse Metso

Seppo Ihalainen Jouni Karvo Ossi Taipale

Helsinki University of Technology Department of Computer Science and Engineering Telecommunications Software and Multimedia Laboratory

Teknillinen korkeakoulu Tietotekniikan osasto Tietoliikenneohjelmistojen ja multimedian laboratorio Distribution: Helsinki University of Technology Telecommunications Software and Multimedia Laboratory P.O.Box 5400 FIN-02015 HUT Tel. +358-9-451 2870 Fax. +358-9-451 5014

\begin{matrix}c \\ ◯ \end{matrix}1999 IPMAN project

CONTENTS

Contents 1

1 Network management 3

1.1 Configuration management . . . . . . . . . . 3

TCP/IP networks . . . . . . . . . . . . . . . . 4 1.2 Security management . . . . . . . . . . . . . . 4

Authentication and authorization . . . . . . . 6 Protocols and programs for network security . . 8 Simple Network Management Protocol (SNMP) 9 Access control tools . . . . . . . . . . . . . . . 10 Security problems of Internet . . . . . . . . . . 11

1.3 Fault management . . . . . . . . . . . . . . . 12 Transferring information . . . . . . . . . . . . 13 Troubleshooting and fault localization . . . . . 14 Testing . . . . . . . . . . . . . . . . . . . . . . 15

1.4 Accounting management . . . . . . . . . . . . 16 Internet pricing . . . . . . . . . . . . . . . . . 17

Bibliography 19

2  IP NETWORK MANAGEMENT

1 NETWORK MANAGEMENT

This chapter describes network management. Section 7.1 describes network configuration management. Section 7.2 studies security management and elements, that are used to secure risky areas of networks such as, hubs, routers, and firewalls. Section 7.3 studies fault management such as troubleshooting, fault localization and testing methods. Section 7.4 studies accounting management.

1.1 Configuration management

Configuration management means initializing and shutting down parts of the network (for example routers, hubs, repeaters) and reporting the changes. It is also concerned with maintaining, adding and updating the relationships among components and the status of components during network operation. Network traffic patterns and identified bottlenecks, that reduce performance, must be understood. Nowadays modern components and subsystems can be configured to support many different applications. The same device can be configured to act either as a router or as an end system node or both. When it is decided, how a device is to be used, the appropriate software and set of attributes and values are chosen for that device [21, page 481]. Reconfiguration is necessary when the network is expanded or in the fault isolation.

As the network scales up in physical size (capability and complexity), so the management capabilities must be enlarged. The aim is that these actions can be automated. It should also be possible to make on-line changes without affecting the entire element or network [17]. Large-scale network management systems must be constructed to support several network elements. They must also be extensible to support new elements. They must be flexible enough to support the rapid deployment of new highly customized services.

Activities in the network management are [7]:

• Activities, that do not affect the function of the element (for example changing the parameters of a device),

• Activities, that affect the function of the element (for example switching off an element in a subnetwork), and • Activities, that make elements to do desired activities

(for example restarting an element).

Dynamic updating of the configuration needs to be done periodically to ensure that the existing configuration is known. This is essential in fault management as well. Configuration management is well organized when all the gathered information and operations are in statistical form. Configuration management tools have reporting components. When network configuration changes, users must be informed about new network elements and resources.

There is a risk of spending money on hardware and services that remain underutilized. On the other hand the risk of underprovisioning is that there is loss of productivity which reflects to the service level [5]. When the network is designed, it is essential to predict the growth of the network. The network should also be prepared to the varying volumes of users.

By continuously addressing the cost of maintenance (Mean Time Between Failure (MTBF) and Mean Time To Repair (MTTR) statistics, costs associated with maintaining) the network as a system can be tuned [23].

TCP/IP networks

TCP/IP networks require more work for system administrators than other networking systems. Administrators have to manually configure each computer for network use when it is added to the network, or when it is moved from one subnet to another. Each computer must manually be assigned a unique IP address and various configuration parameters must be set.

There is a need for tools, that automatically assign addresses and set configuration parameters. Some client/server solution are already available. Client machines find the details of other hosts on the network using the Domain Naming Systems (DNS) protocols, and can be told their network configuration using Bootstrap (BOOTP) protocol. Static allocation of IP addresses to particular machines can be chosen over dynamic allocation using the Dynamic Host Configuration Protocol (DHCP) [1].

1.2 Security management

Security management covers such areas as detecting, tracking and reporting security violations, and creating, deleting and maintaining security-related services such as encryption, key management, and access control. Distributing passwords and secret keys to bring up systems is also a function of security management [25, page 12].

As computer-based communications and networks that link open systems continue to expand, security management becomes critical. Nevertheless, standardization of security properties has developed slowly. Security of devices and networks must be compared to possible threats and risks. If the risks are high, the devices and the networks must be provided with more reliable secure properties.

Basic security services, that are defined in ITU-T (International Telecommunication Union - Telecommunication standards) Recommendation X.800, are:

• Access control The property of controlling network and computer resources in such way that only legitimate users can access them within their limits. One approach is to attach to an object a list which explicitly contains the identity of all permitted users (an Access Control List (ACL)) [20]. Access control tools are discussed in subsection 7.2.X.

• Authentication The property of knowing that the data received is the same as the data that was sent and that the claimed sender is in fact the actual sender. Several authentication techniques have been developed for example, technologies that provide passwords that are only used once (commonly called one-time passwords) and Kerberos. Kerberos is a software designed and developed at Massachusetts Institute of Technology (MIT) to perform distributed authentication in an open network environment.

• Confidentiality The property of communicating such that the intended recipients know what was being sent but unintended parties cannot determine what was sent. Encryption is commonly used to provide confidentiality.

• Integrity The property of ensuring that data is transmitted from source to destination without undetected alteration.

One way to provide integrity is to produce a checksum of the unaltered file, store that checksum offline, and periodically (or when desired) check to make sure the checksum of the online file hasn’t changed (which would indicate the data has been modified) [20].

• Non-repudiation

The property of a receiver being able to prove that the sender of some data did in fact send the data even though the sender might later desire to deny ever having sent that data [3].

Effective security management must be involved in all steps of the data storage and transfer process. Logs are important security tool and therefore security management is involved with the collection, storage and examination of the audit records and security logs. Increasing the level of network security will affect to the openness of the system and to the cost of maintaining the network. The implementation of security management with proper controls should not affect the normal operation of network personnel. It is problematic to decide where to locate data security properties.

Authentication and authorization

Almost all applications utilize user information and presume an authentication of users. Authorization is the determining whether an identity is permitted to perform some action, such as accessing a resource [12]. Passwords, smart cards and certificates are used to authenticate a user. A user may have the rights to use more than one name and a user may have rights to use identities established by multiple organizations (such as universities and scholarly societies). There might be an advantage if all the user information is available in same directory. All the applications could then use the same information. Users have to log in only once to be able to use all the services and resources [13].

There are some basic requirements for authentication [12]:

• access management solution needs to work at a practical level

• the solution needs to be secure • it should make access easier, minimizing redundant au-

thentication interactions and providing user-friendly information resources

• it needs to scale

• it needs to be robust

• there should be a simple and well-defined (standard) interface between resource operator and licensing institution

• it must be able to recognize the need for a user to access a resource independent of his or her physical location (for example, a user must be able to connect to the internet via a commercial ISP, a mobile IP link, or a cable television internet connection from home)

The basic access management problem is licensing agreements for networked information resources. The situations where institutions agree to share limited access are difficult. At present, most access to network information resources is not controlled on a fine-grained basis. There is a danger that by accommodating all the needs for fine-grained access management into the basic access management mechanisms will produce a too complex and costly system [12].

Management data is a problem in the current access framework. The problem is the conflict between private and public data. Most of the data have to be sorted out at the institutional policy level and it may involve making sacrifices in order to ensure privacy. Some institutions may be legally limited in their ability to collect certain management data.

Proxies and credential-based authentication (the user presents a credential to the operator as evidence that he or she is a member of the user community) schemes seem to be viable. Proxy servers will become a focal point for policy debates about privacy, accountability and the collection of management information. Successful operation of a proxy server means that the user trusts the licensee institution to behave responsibly and to respect privacy.

A cross-organizational authentication system based on a credential approach has the advantage of greater transparency. Resource operators can have a higher level of confidence in the access management mechanisms and a greater ability to monitor irregular access patterns. Privacy, accountability and collection of management statistics must be taken up for discussion among a larger group of parties.

IP source filtering does not seem to be a viable solution, although it may be very useful for some niche applications, such as supporting public workstations. It can be used more widely, but it cannot support remote users flexibly in its basic form. Most real-world access management systems are going to have to employ multiple approaches and IP source address filtering is likely to be one of them [12].

Protocols and programs for network security

Table 1.1 describes protocols and programs, that are developed to provide security over the Internet. Secure-HTTP (S-HTTP) is an application-level protocol that provides security services across the Internet. It provides confidentiality, authenticity, integrity, and non-repudiability. S-HTTP is limited to the specific software that is implementing it, and it encrypts each message individually [2].

Level  Protection Used
Application-level  S-HTTP, SSH, stelnet, S/MIME
Transport-level (TCP, UDP)  SSL, PCT
Network-level (IP)  IPSEC

Table 1.1: Protocols and programs developed to provide security over the Internet

Secure Shell (SSH) and Secure telnet (stelnet) are programs that allow you to log in to remote systems and have a encrypted connection. SSH uses public-key cryptography to encrypt communications between two hosts, as well as for user authentication.

Secure Multipurpose Internet Mail Extension (S/MIME) is an encryption standard used to encrypt electronic mail, or other types of messages on the Internet. It is an open standard developed by RSA Data Security Inc.

Secure Socket Layer (SSL) is an encryption method developed by Netscape to provide security over the Internet. SSL is a protocol layer that is located between the network layer and applications, so that, in theory, it can be used with any application. However, it is vulnerable to poor application design. It provides data encryption, server authentication, message integrity, and optional client authentication for a TCP/IP connection [2]. The Private Communication Technology (PCT) protocol is developed by the Microsoft company to be used mainly in their Internet Explorer browser.

IP Security Architecture (IPSEC) provides a standard security mechanism and services to the currently used IP version 4 (IPv4) and to the new IP version 6 (IPv6). IPSEC is less dependent of individual applications than SSL. It provides IP-level encryption by specifying two standard headers: IP Authentication Header (AH) and IP Encapsulating Security Payload (ESP). IP Authentication Header provides strong integrity and authentication. It computes a cryptographic authentication function over the IP datagram and use a secret authentication key in the computation. IP Encapsulating Security Payload provides integrity and confidentiality to IP datagrams. It encrypts data to be protected and places the encrypted data in the data portion of the IP Encapsulating Security Payload. These mechanisms do not provide security against traffic analysis and other methods should be used to protect against it. Also any specific protocol for key management are not provided, only requirements for such systems to be used in conjunction are described [16].

Simple Network Management Protocol (SNMP)

The SNMP architecture makes a distinction between message security services (integrity, authentication and encryption) and access control services. Both, message security and access control services can be provided by multiple security or access control models. The architecture allows the coexistence of multiple models in order to allow future updates, in case the security requirements change or cryptographic protocols need to be replaced [19, pages 690–692].

The User-based Security Model (USM) provides integrity, authentication and privacy, and is the standard security model currently used with SNMP version 3 (SNMPv3). The View-based Access Control Model (VACM) provides checking whether the users have proper access rights to access one or more objects in a Management Information Base (MIB) and perform operations on these objects. USM is discussed in RFC 2274 and VACM in RFC 2275.

Access control tools

Hubs, routers and firewalls are used to limit the access to the networks or the parts of the networks.

Hubs used in LANs are provided with simple security properties for example, tapping or listening protection, penetration prevention. The security level operations can be restricted by the date or by the time period.

Routers are provided with security properties as well. The network security is improved by routers, but they bring along new problems. For example, routing protocols are susceptible to security attacks, and routing mistakes may allow the entrance of unauthorized personnel into the network. The unlicensed remote and local operations must be prevented with a username. The traffic control through the router is managed with a packet filtration which is based on the list. Some properties can be attached to the router, such as the network address translation (NAT), port address translation (PAT), logging of events and alerts. It has not been possible to expand the capacity of traditional routers as cost-effectively as the capacity of PC workstations or the network traffic volume [10].

Firewall systems control connections between the company’s system and the outside world. There are two main types of firewalls: filtering and proxy. A filtering firewall routes packets based on the addresses and the types of the packets. A proxy firewall controls the traffic at the application level.

The firewall is often the bottleneck of the network. Examining every packet before transfer requires an efficient processor. Also, the configuration failures may leave unprotected passages through the firewall.

The firewall can be implemented in various ways:

• The screened subnet architecture,

• The multi-homed host architecture, and

• The router with security property, also called firewalls.

The screened subnet architecture usually consists of interface router, open network, security server (called Bastion Host) and internal router. Interface router is the only connection to outside from organization’s own network. It checks that packets coming from outside don’t have source address of the internal network. The internal router filters packets by address, protocols and port numbers. It controls services available from both outside and inside of the organization’s network. Bastion Host is the security server. Security server is the connection point with limited services in the intranet and in the extranet e.g. Domain Name Service (DNS), Simple Mail Transfer Protocol (SMTP) and File Transfer Protocol (FTP) [9, pages 398–399].

The multi-homed host architecture is a solution where everything is in a single host. In multi-homed host all services needed by firewall are carried out in one host. IP-router is replaced by a multi-homed host. The multi-homed host connects the inside and outside network. The multi-homed host doesn’t route packets on IP level but instead transfers packets on the application level [9, pages 399–400].

Security problems of Internet

The list below describes the security problems of Internet, that are current today.

• Weak authentication

Passwords on the Internet can be cracked by a number of different ways. The two most common methods are cracking the encrypted form of the password and by monitoring communications channels of password packets.

Another problem with authentication results from some TCP or UDP services being able to authenticate only to the granularity of host addresses and not to specific users. For example, an NFS (UDP) server cannot grant access to a specific user on a host, it must grant access to the entire host. The administrator of a server may trust a specific user on a host and wish to grant access to that user, but the administrator has no control over other users on that host and is thus forced to grant access to all users (or grant no access at all).

• Ease of spying/monitoring

When a user connects to her account on a remote host using TELNET or FTP, the user’s password travels across the Internet unencrypted, or in plain text. Another method for breaking into systems is to monitor connections of IP packets bearing a username and password, and then using them on the system to login normally.

If the captured password is to an administrator account, then the job of obtaining privileged access is made much easier.

Electronic mail, as well as the contents of TELNET and FTP sessions, can be monitored and used to learn information about a site and its business transactions. Most users do not encrypt e-mail, yet many assume that e-mail is secure and thus safe for transmitting sensitive information.

The X Window System is an increasingly popular service that is also vulnerable to spying and monitoring. X Window permits multiple windows to be opened at a workstation, along with display of graphics and multi-media applications (for example, the WWW browser Mosaic). Intruders can sometimes open windows on other systems and read keystrokes that can contain passwords or sensitive information.

• Host-based security does not scale

Host-based security does not scale well: as the number of hosts at a site increases, the ability to ensure that security is at a high level for each host decreases. Secure management of just one system can be demanding, managing many such systems could easily result in mistakes and omissions. A contributing factor is that the role of system management is often short-changed and performed in haste. As a result, some systems will be less secure than other systems, and these systems could be the weak links that will break the overall security chain [26].

1.3 Fault management

Physical network problems account more than half of all network problems. Locating the origins of such problems as fiber cut, the earth is incorrect, adapter is broken or incorrectly joined, costs network providers time and money. Network management systems have also traditionally focused on the logical connection between the end user and the network destination, since most network problems have been a result of errors created by software applications [17]. Finding and repairing failures of software applications is usually difficult, in such a case that the workstation sends but does not receive packets, too many collisions occur or frames are too short or too long.

Fault management is troubleshooting, fault localization, isolation and correction. Today, the process of fault management needs to be automated. Rapid and accurate correction of network problems need to be created by the network or by the end user. Even if the entire network were violated, the network management application should work [18].

Expert systems provide an efficient and cost effective way of automating network fault management. By automating fault management, problems and possible diagnosis can be done faster and more efficiently. Problematic for rule-based expert systems is the real time performance [6]. By expert systems bounds on response times are difficult to establish. There are also risks pushing the limits of automation through the introduction of new or only partially proven technologies.

Transferring information

There are two mechanisms of transferring network management information from a managed entity to the manager: polling and sending alerts. Alerts are information delivered to a manager that is not the result of a specific query. Advantages and disadvantages exist within each method. Most network management systems use an optimal combination of alerts and polling. Doing so maintains the advantages of each and eliminates the disadvantages of pure polling [22].

Most systems poll the managed objects, search for error conditions and illustrate the problem in a graphic format or a textual message. Disadvantages of polling are response time for problem detection and volume of network management traffic. Having to poll many management information base (MIB) variables per machine on a large number of machines is itself a problem. The ability to monitor such a system is limited. Polling many objects on many machines increases the amount of network management traffic flowing across the network. It is possible to minimize this through the use of hierarchies (polling a machine for a general status of all the machines it polls). Anyway, the response time will be a problem [22].

If a system fail shortly after being polled, there may be a significant delay before it is polled again. During this time, the manager must assume that a failing system is acceptable.

While improving the mean time to detect failures, it might not greatly improve the time to correct the failure. The problem will generally not be repaired until it is detected.

The problems of sending alerts are potential loss of critical information and possibility to over-inform the manager. An ideal management system will generate alerts to notify its management station of error conditions. Alerts are usually not to be delivered when the managed entity fails or the network experiences problems. It is important to remember that failing machines and networks can not be trusted to inform a manager that they are failing. The manager should periodically poll to ensure connectivity to remote stations, and to get copies of alerts that were not delivered by the network [22].

Alerts in a failing system can be generated so rapidly that they impact functioning resources. An ”open loop” system in which the flow of alerts to a manager is fully asynchronous can result in an excess of alerts being delivered. There may be a situation where all available network bandwidth into the manager is saturated with incoming alerts and this could prevent the manager from disabling the mechanism generating the alerts. Methods are needed to limit the volume of alert transmission and to assist in delivering a minimum amount of information to a manager. Alarm correlation is done by filtering secondary alarms. Expert systems are used in filtering alarms.

Many management tools also log events with different formats and different sources. These events should later be correlated using time stamp to identify the source of the problem. Also topology information is needed to identify the precise location of the problem.

Many devices have buffers reserved for logs. The buffer takes disk space from the other operations. When the buffer is full, new logs are overwritten to the oldest ones. If the buffer becomes full quickly, some important logs may disappear before they are noticed.

Troubleshooting and fault localization

Several diagnostic tools are used for troubleshooting and fault localization purposes. These are for example, ifconfig, arp, netstat, ping, nslookup, dig, ripquery, traceroute, and etherfind [8, pages 261–262].

The packet inter net groper (ping) tests whether a remote host can be reached. It sends an internet control message protocol (ICMP) loopback request packet to a target IP address. The receiver accepts the loopback request and issues a loopback response to the initiator of the Ping command. It cannot be determined with echo-response which one of the possible routes is violated, because loopback request and loopback response may use different routes. Ping can be used only to determine if the target IP-address is available. Ping can be used to measure response time of different packet sizes as well [5, page 33].

Unlike ping traceroute forces every router to send back an ICMP-control message. Most Traceroute applications send a sequence of User Datagram Protocol (UDP)-package to a UDP-port, which is randomly selected [4]. Sometimes firewalls filter these packets from the main traffic, when the tracing has come to an end.

After detecting a fault, operators should as soon as possible find the root cause of the problem. The root problem is a network fault, which the network manager should investigate and repair, such as fiber cut, an equipment failure or a bandwidth overload. It is important to clearly define the problem so that it describes the root problem, not a symptom [15]. Operators must also determine what troubles it causes and to whom.

At first it is useful to find the parts of the network, where the fault can not be, by eliminating processes. These parts can be neglected. If there is a very serious and difficult problem, the network can be divided into smaller parts and the parts are inspected one by one. Also connectivity tests is used to find the violated devices. By the same time can be detected if the fault is local or widespread. By testing the data integrity can be found out whether a part of the packets are lost by transmission. Delays are also tested, because some faults are results of excessive delays [4].

Testing

Measuring the accessibility of devices helps us to collect information about the status of the network. In general, some points are selected to measure connections. Measurements are done by a control device. After measurements, the results shall be interpreted correctly. Measuring the accessibility is not a complete method. For example, controlling only routers is not enough. If the setup of the router has been specified so that it gives the preference to packet switching, it will send packets normally although it is heavily loaded and it perhaps can not answer to messages of controlling device. Another problem is for example that two devices can be reached but they do not have connection with each other.

Another method is to control routing tables. Routing should change only when the topology of the network changes. If there are changes at other times, there is probably something wrong with the network. Routing tables express the topology of the network. The program of finding out the network topology from routing tables is more difficult than the program of controlling accessibility [4]. In order to get information from routers to the routing tables, routers must be Simple network management protocol (SNMP) compatible, dynamic routing tables must support sending request or routing information must be tapped from the routing protocol. These requirements are not necessarily realized although SNMP is common. Analysis of the routing tables can concentrate on the parts of the network which are most susceptible to failures.

1.4 Accounting management

Accounting management includes mechanisms to monitor and control information and resources which concern individual users. The accounting management includes:

• Usage measurement Data collection for charging and processing of data has to be reliable and, sometimes, has to be done in real time.

• Tariffing/pricing A tariff is a set of data used to determine the charges for services used. It depends on the service, origination and destination, tariff period, and day.

• Collections and finance This includes administration of customer accounts, informing customers an payment dates, payment amount, and collection of payments.

• Enterprise control Enterprise control is responsible for the proper financial management of an enterprise. It includes identifying and ensuring financial accountability of officers. Also checks and balances needed for financial operation of an enterprise are included [25, pages 64–66].

A system which generates data for accounting purposes is called an accounting management agent. Accounting managers are systems which interrogate accounting management data or obtain it in other ways. If the accounting management is distributed across various systems, all systems may be required to control their own area themselves. Furthermore, a system may request information from other systems in order to square its accounts [11, pages 188–189].

Accounting data is sensitive. The collector must provide confidentiality for accounting data at the point of collection, through transmission and up to the point where the data is delivered. The delivery function may also require authentication of the origin and destination and provision for connection integrity (if connections are utilized). Security services can be provided for example, by SNMP and its mechanisms.

Internet pricing

Components of Internet pricing include three basic elements (see figure 1.1). An access-fee is collected which is usually a monthly charge for using an access link to the network. The price depends on the capacity of that link. The setting up connections or reservations can be charged separately. Usage-fee can be used to charge services on time-, volume-, or QoS-basis. This fee determines the actual resource usage customers consume. Content-fee depends on the application content. It may be omitted (e.g., telephony, fax, e-mail services where the content is provided by the customer herself), billed separately (e.g., Helsingin Sanomat on-line edition), or integrated into the telecommunication charging system (e.g., 0900 numbers) [24].

Figure 1.1: Components of Internet pricing [24]

The current pricing model is based on the assumption of a single service best-effort network that provides a similar service to all customers. Seller and customer do not have a direct control over the actual service, e.g., in terms of parameter determining volume, connection time, and quality. For example, current Internet service, except for the dialup access, does not allow billing based on time because the IP is a connection less communication, and timing information about the beginning and ending of billing is too difficult to obtain.

Accounting is usually based on mechanisms offered by commercially available routers and switches. The most often used approach employs packet filtering and statistical sampling. However, it is difficult to account usage-based traffic even single users, since the granularity of these methods is too coarse and the measurement overhead significant [24].

A final data collection concern of routers is whether packets should be counted on entry into a router or upon exit from a router. The nature of IP is, that not every packet received by a router is actually passed to an output port. The Internet Protocol allows routers to discard packets for example, in times of congestion when the router cannot handle the offered load. It is presumed that higher level protocols (e.g., TCP, UDP) will provide delivery service the user seems necessary. Another question is, whether an internet accounting system should count all packets offered to a router or just those that are finally passed by the router to a network (should a user pay for undelivered packets) [14].

For volume measurements the IETF Real-time Traffic Flow Measurement (rtfm) working group has proposed standards to meter flows and to distribute this accounting information via SNMP [24].

The Remote Authentication Dial-In User Service (RA-DIUS) is a protocol specified by the Internet Engineering Task Force (IETF) working group. It helps managing Internet access networks. Since these links are sensitive to security and accounting, a protocol is provided to authenticate dial-in users and negotiate configuration data. RADIUS servers are implemented by most router manufactures. Accounting data can be collected on a time-, packet-, or octet-basis for a particular service [24].

BIBLIOGRAPHY

[1] Alexander, S., and Droms, R. DHCP options and BOOTP vendor extensions. RFC 2132 (1997).

[2] Arnaud, D. Security status and issues & electronic commerce on the internet, July 1995. August 18, 1999, http://ecwww.eurecom.fr/˜ arnaud/zds/report/report.html.

[3] Atkinson, R. Security architecture for the internet protocol. RFC 1825 (1995).

[4] Ballew, S. M. IP-verkkojen hallinta Ciscon reittimillä. Suomen Atk-kustannus Oy, Helsinki, Finland, 1998.

[5] Blommers, J. Practical Planning for Network Growth. Prentice Hall PTR, New Jersey, USA, 1996.

[6] Fuller, W. Network management using expert diagnostics, 1999. April 26, 1999, Summit On Line, White Papers, http://www.summitonline.com/netmanage/papers/stanford1.html.

[7] Hautaniemi, M. Tkk/atk-keskuksen tcp/ip-vekon valvonta ja hallinta. Master’s thesis, Department of Computer Science and Engineering, Helsinki University of technology, 1994. April 25, 1999, http://www.hut.fi/˜ hau/thesis/verkonhall toteutus.html.

[8] Hunt, C. TCP/IP Network Administration. O’Reilly & Associates Inc., 1992. California, USA.

[9] Hunt, C. TCI/IP Verkonhallinta. Suomen atkkustannus, Helsinki, Finland, 1998.

[10] ICL. Icl:n verkkoaapinen, verkkoratkaisut ja palvelut, 1999. August 25, 1999, http://www.icl.fi/.

[11] Kauffels, F.-J. Network Management, Problems, Standards and Strategies. Addison-Wesley Publishing Company, New York, USA, 1992.

[12] Lynch, C. A white paper on authentication and access management issues in cross-organizational use of networked information resources, 1998. May 5, 1999, Coalition for Networked Information Revised Discussion Draft, http://www.cni.org/projects/authentication/authenticationwp.html.

[13] Mensola, S. Ip-verkon kommunikaatiopalveluiden hallinta. Master’s thesis, Department of Electrical and Communications Engineering, Helsinki University of Technology, 1998. May 3, 1999, http://kyyppari.hkkk.fi/˜ k23332/dippa/luku2.htm.

[14] Mills, C., Hirsh, D., and Ruth, G. Internet accounting: Background. RFC 1272 (1991).

[15] Network General Corporation. Proactive solutions to the five most critical networking problems, 1997. April 21, 1999, Summit On Line, http://summitonline.com/netmanage/papers/netgen2.html.

[16] Pulkki, A. The ip security architecture. In Proceedings of the HUT Network Seminar ’96 (1995), Department of Physics, Helsinki University of Technology. July 4, 1999, http://www.tcm.hut.fi/Opinnot/Tik-110.501/1995/ip-sec-arch.html.

[17] Robinson, C. Integrated network management for multimedia networking, 1999. April 19, 1999, http://engineer.home.mindspring.com/book.htm.

[18] Rose, M. T. The Simple Book, An Introduction to Management TCP/IP-based internets. Prentice Hall, USA, 1991.

[19] Schönwälder, J., and Quittek, J. Secure management by delegation within the internet management framework. In Integrated Network Management VI (Boston, USA, May 1999), pp. 690–692.

[20] Site security handbook. RFC 2196 (1997). Fraser, B., Ed.

[21] Stallings, W. Local and Metropolitan Area Networks, 4 ed. Maxwell MacMillan International, New York, USA, 1993.

[22] Steinberg, L. Techniques for managing asynchronously generated alerts. RFC 1224 (1991).

[23] Stevenson, D. W. Network management - what it is and what it isn’t, 1995. May 26, 1999, http://netman.cit.buffalo.edu/Doc/Dstevenson.

[24] Stiller, B., Fankhauser, G., Plattner, B., and Weiler, N. Charging and accounting for integrated internet services - state of the art, problems, and trends. In The Internet Summit, INET’98 (Switzerland, July 1998).

[25] Udupa, D. K. TMN Telecommunications Management Network. McGraw-Hill, New York, USA, 1999.

[26] Wack, J. P., and Carnahan, L. J. Keeping Your Site Comfortably Secure: An Introduction to Internet Firewalls. NIST Special Publication 800-10, U.S.Department of Commerce, National Institute of Standards and Technology, 1999. August 15, 1999, http://csrc.nist.gov/nistpubs/800-10/main.html.